G(t)=9-2t, write G(-3+h) - G(-3) in Simplest form.
Answer by longjonsilver(2297)

if g(t) = 9-2t, then

g(-3+h) = 9-2(-3+h) --> can you see how obvious this line is?
g(-3+h) = 9+6-2h
g(-3+h) = 15-2h

and g(-3) = 9-2(-3)
g(-3) = 9+6
g(-3) = 15

So, g(-3+h) - g(-3) = 15-2h - 15
g(-3+h) - g(-3) = -2h

jon.